• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name LeRose

    Log entry time 00:00:49 on July 08, 2010

    Entry number 330724

    keyword=Shift Summary

    Day shift 7/7/2010
    John LeRose (SL)
    James Beacham (TO)
    Sophia Iqbal  (3rd)
    
    At shift start the accelerator is down due to a weather induced
    power shut down.
    
    The magnets, except for the right dipole, have been powered down.
    Expect to return to full power at 20:00
    
    19:07 Power reduction officially ends
    19:26 We get 1st word (e-mail) that power reduction has ended
    	Shortly after we get the word to start repowering the spectrometers
    20:00 Spectrometers back at 1.131 GeV/c (Q2/3's cycled to 1000 Amps)
    	Wait for beam.
    21:21 Still waiting
    20:00 MCC says at least another hour to beam
    23:50 MCC says they're about to send us tune beam
    22:58 MCC says tune beam looks OK. Going to CW raster off to check the beam 
    	spot on the beam dump